The length of a rectangle is 2 cm more than 5 times its width. If the area of the rectangle is 65 cm2, find the dimensions of the rectangle to the nearest thousandth.
Hint: Call the width x. Then the length is 5x + 2. Now write your equation and solve.

Answer by THANApHD(104)

Call the width x. Then the length is 5x + 2.
Area = length x width
Area = x(5x+2)
65 = x(5x+2)
65-65= 5x^2+2x-65
0= 5x^2+2x-65
Solve for x using quadratic formulae